OpenStudy (anonymous):

Could somebody please give me a hand with this question? Find the line tangent to the curve y=x^2-3x+2 and parallel to the line 4x-2y=7.

OpenStudy (campbell_st):

ok... so can you differentiate \[y = x^2 - 3x + 2\]

OpenStudy (anonymous):

y=2x-3

OpenStudy (campbell_st):

ok... thats a good start. next whats the slope of the given line..?

OpenStudy (anonymous):

m=2/1

OpenStudy (campbell_st):

ok m = 2 so you now need to equate the derivative and the slope of the line to find the x value of the point where the tangent is parallel so solve 2 = 2x - 3

OpenStudy (anonymous):

x=5/2?

OpenStudy (campbell_st):

ok... now you need to find the point on the curve where the slope is 2 substitute x = 5/2 into your original curve to find the y value for the ordered pair.

OpenStudy (anonymous):

Into y=x^2-3x+2?

OpenStudy (campbell_st):

thats it ... it will be a fractional answer

OpenStudy (anonymous):

Into the derivative of the curve or the original function?

OpenStudy (campbell_st):

the original...

OpenStudy (anonymous):

*sigh* So simple, yet so difficult..... Thanks for the help, I really appreciate it.

OpenStudy (campbell_st):

lol... yep... one of the keys is realizing that the 1st derivative of an equation gives the equation of the slope of the tangent,. so you line is y = 2x + b substitute x = 5/2 and y = to get your line.. hope it helps